OpenStudy (anonymous):

Write the equation of the circle with center (0, 6) and radius 1.

OpenStudy (anonymous):

x^2 + (y-6)^2 = 1

OpenStudy (dumbcow):

(x-h)^2 + (y-k)^2 = r^2 center is (h,k) r is radius just substitute the numbers into general formula